36 marbles colored either red or blue. Twice the number of red marbles is 6 less than the number of blue marbles. How many red marbles are in the bag?
Answer by TimothyLamb(4379)

x = number of red marbles
y = number of blue marbles
---
x + y = 36
2x = y - 6
---
put the system of linear equations into standard form
---
x + y = 36
2x - y = -6
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = number of red marbles = 10
y = number of blue marbles = 26
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---